OpenStudy (anonymous):

Write the polynomial of the smallest degree with roots 8, 2i, and -2i

OpenStudy (anonymous):

(x-8) (x-2i) Right?

OpenStudy (anonymous):

mhmm

OpenStudy (amistre64):

is a poly defined with complex coeffs? or just real coeefs?

OpenStudy (anonymous):

(x-8) (x-2i) (x+2i)

OpenStudy (anonymous):

expand that

OpenStudy (anonymous):

x^3-8*x^2+4*x-32

OpenStudy (amistre64):

hmmm, the definition of a poly seems to only deal with the variables and not the coeffs; there are complex polynomials - or, polys with complex coeffs

OpenStudy (anonymous):

you are thinking too much @amistre64

OpenStudy (amistre64):

i agree that my conundrum has no real bearing in the question, but it does seem to be an oddity for me :)

OpenStudy (anonymous):

Her q is Write the polynomial of the smallest degree with roots 8, 2i, and -2i there are 3 roots so just fill in the blanks of this form (x-a) (x-b) (x-c) what a b and c is 8, 2i, and -2i That simple then just expand

OpenStudy (amistre64):

ah, i missed the last root on the end i forgo the swapping signs and just subtract x from each rot, then expand 8, 2i, -2i 8-x, 2i-x, -2i-x f(x) = (8-x)(2i-x)(-2i-x)

OpenStudy (anonymous):

lol well roots of imaginaries always come in pairs right? :D

OpenStudy (anonymous):

(x-8)(x^2+4) Then expand. (X^2+4) is the imaginary roots. Imaginary roots always take this form (x^2 + b^2) where bi is a root

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%28x-8%29%28x-2i%29 i think pairs of complex only exist with real polys; this construct only has one complex root

OpenStudy (shubhamsrg):

you may also use sum of roots,product in pairs and product of roots alternately..

OpenStudy (shubhamsrg):

alternatively*

OpenStudy (anonymous):

8, 2i, -2i 8-x, 2i-x, -2i-x f(x) = (8-x)(2i-x)(-2i-x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that looks wrong or just harder mines is right (x-8) (x-2i) (x+2i) Expand x^3-8*x^2+4*x-32